Calculate the molarity of a solution having a concentration of 246.00 micrograms/all CaCO3
1
Expert's answer
2020-10-05T14:26:39-0400

Molarity is a unit of concentration, measuring the number of moles of a solute per liter of solution.

n(CaCO3) = m/M = 246/100 = 2.46 micromoles

Molarity = n/V = 2.46 micromoles/1 L solution